# How Does A Text Preprocessing Pipeline Affect Syntactic Ontology Matching?

### v4

@@ 4.2.3. Different Stemming Methods
Figure 6 shows the comparison of different stemming methods: Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL). For SP and SS, all data points located on the equivalent line indicate that there is no difference between SP and SS. For SP/SS and SL, most data points above the equivalent line in precision and F1 score indicate that SP/SS is more effective than SL. In the violin plots: (a) Precision: The median number of SP and SS is greater than that of SL, and there is no difference between SP and SS. The shape of the distribution is identical. (b) Recall: The median number and the shape of the distribution are identical in SP, SS, and SL. (c) F1 Score: The median number of SP and SS is greater than that of SL, and there is no difference between SP and SS. The shape of the distribution is identical in SP, SS, and SL.
@@ 4.2.4. Lemmatisation with/without POS Tagging
Figure 7 shows the comparison of Lemmatisation (L) and Lemmatisation + POS Tagging (LT). All data points located on the equivalent line indicate that using POS Tagging in Lemmatisation does not help syntactic OM. In the violin plots, for each of (a) Precision, (b) Recall, and (c) F1 Score, the median number and the shape of the distribution are identical for L and LT.
@@ 4.3. Discussion
    Phase 1 text preprocessing methods: We observe that matching performance usually increases with each Phase 1 text preprocessing method. However, Phase 1 methods could also benefit from customisation in OM tasks. Some traditional methods originating from natural language processing (NLP) are not useful for OM and need to be adjusted. For example, sentence segmentation (i.e. splitting long text into sentences) is not applicable for ontology entities because they are generally short text fragments and do not need such operations. Word tokenisation (i.e. breaking text into single words) could be rewritten to detect the abbreviations that are common practice in ontology concept names and to tackle the conflicting use of camel case and snake case.
@@
      Stop Words Removal (R): “AND”, “OR”, and “NOT” are stop words in English and usually carry little useful information, whereas these words express logical operations in ontology entities and may carry important semantics.
@@
      Stemming (S) vs. Lemmatisation (L): While L tends to avoid generating FPs, it may also miss some implicit TPs. On the other hand, S is more aggressive in finding TPs, but aggression can also lead to more FPs being found. Based on our experiments with no post hoc correction applied, L is better than S.
@@
      Different stemming methods: Porter Stemmer (PS) and Snowball Stemmer (SS) have been found to perform better than Lancaster Stemmer (LS), and we cannot see a significant difference between Porter and Snowball. Although the Lancaster Stemmer is more aggressive in detecting more TPs, it does not lead to performance improvement as more FPs are matched.
@@ 5.1. Causal Analysis
    Phase 1 text preprocessing methods (Tokenisation and Normalisation) do not change the text meaning. For example, isReviewing is equivalent to is_reviewing. This means that applying Phase 1 methods only helps to detect TPs, while the number of FPs remains unchanged. For this reason, Phase 1 methods always have a positive effect on precision, recall, and overall F1 score.
@@
    Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation) are actually relaxations of matching rules in OM tasks. Moving through the text preprocessing pipeline tends to detect more TPs and FPs in the derived alignment (A). For example, isReviewing and isReviewedBy may be object properties with distinctly different meaning, but removing the common stop words “is” and “by”, and using Stemming/Lemmatisation to retrieve the same root word “Review”, could cause a false match. For this reason, recall is always increasing, but precision and overall F1 score are less reliable.
@@
    If we define $\Delta TP$ and $\Delta FP$ as the increase in TP and FP for a preprocessing method, then the threshold to increase precision and F1 score is:
@@
    In our experiments, we actually observe a reduction in precision and overall F1 score. This means that Phase 2 methods produce more FPs than TPs, and this proportion is less than the original number of $TP/FP$. So performance does not improve, unless the benefit of each TP is considered more valuable than the disbenefit of each FP. This could apply, for example, if we are expecting a post hoc correction phase where removing FPs is considered to be an easier human task than adding missing TPs.
@@ 5.2. Repair Method
While Phase 2 text preprocessing pipeline methods have been shown to be less effective in syntactic OM mainly caused by FPs, we propose a pipeline repair approach to differentiate FPs from TPs. As illustrated in Figure 8: (1) Phase 1 methods shift Alignment (A) towards Reference (R). The number of TPs increases, while the number of FPs decreases. (2) Phase 2 methods expand Alignment (A). The number of TPs increases, but the number of FPs also increases. (3) Our context-based pipeline repair aims to collapse Alignment (A). The number of FPs significantly decreases, with a slight decrease in TPs.
@@
One critical step in our approach is to retrieve a reserved word set that may cause FPs after the text preprocessing, and these words will be excluded before the text preprocessing. The selection criteria are based on two widely agreed assumptions: (1) there are no duplicate entities within a single ontology; and (2) ontologies that represent the same domain knowledge tend to use similar terminologies (we call our approach context-based here because pairs of words may have the same or different meanings in different contexts). Based on these two assumptions, we propose a simple and intuitive Algorithm 1 to retrieve the reserved word set for context-based pipeline repair. For both $O_{s}$ and $O_{t}$, the algorithm first iterates on all pairs of entities $e_{i},e_{j}$ in each of them. For a specific text preprocessing method $f(\cdot)$, if $f(e_{i})=f(e_{j})$, we retrieve the different words between $e_{i}$ and $e_{j}$ and store them in the reserved word set. If a word appears in the reserved set, the text preprocessing pipeline skips the operation for this word. To simplify the reserved word set, we also remove the words where $f(w)=w$ from the final set because either skipping or keeping these words in the reserved word set would not change the mapping results.
@@
  /* Phase 2: Find duplicates in $Reserved\_Word\_Set$/
@@
An example of generating and using a simple reserved word set is illustrated below. (1) The two object properties was_a_member_of and has_members from a single ontology have the same result “member” via traditional text preprocessing. Because there are no duplicate entities within a single ontology, we use a reserved word set in our proposed pipeline repair approach to determine that they are distinguishing entities. The initial step (i.e. Phase 1 of Algorithm 1) is to add [“was”, “a”, “member”, “of”, “has”, “members”] to the reserved word set so that these two object properties would not have the same text preprocessing results. was_a_member_of preprocessed with skipping the reserved word set is “was a member of”, while has_members preprocessed with the reserved word set is “has members”. The revision step (i.e. Phase 2 of Algorithm 1) is to check whether there are duplicates in the reserved set. We can observe that the word “member” is a duplicate because it is the same before and after text preprocessing. Removing this word from the reserved word set still makes the two object properties different. Therefore, the final reserved word set is [“was”, “a”, “of”, “has”, “members”]. (2) The generated reserved word set can be used to repair false mappings between entities within the same domain context but coming from different ontologies. For example, we expect that the two object properties has_a_steering_committee and was_a_steering_committee_of are non-identical. While a false mapping may occur when they both have the same result “steer committe” after the traditional text preprocessing, using the reserved word set can repair this false mapping. As the words “has”, “a”, “was”, and “of” are listed as reserved words, these two named properties are processed as “has a steer committe” and “was a steer committe of”, respectively.
@@ 5.3. Evaluation
We apply our context-based pipeline repair approach to the same OAEI datasets as above. Figure 9 compares with and without context-based pipeline repair in 8 tracks with 49 distinct alignments. The text preprocessing pipeline methods implemented include all Phase 2 text preprocessing methods: Stop Words Removal (R), Porter Stemmer (SP), Snowball Stemmer (SS), Lancaster Stemmer (SL), Lemmatisation (L), and Lemmatisation + POS Tagging (LT). (a) Precision: Most data points are above the equivalent line, indicating that context-based pipeline repair significantly improves the matching correctness. (b) Recall: Most data points are located on the equivalent line, and only a few data points are below the equivalent line, indicating that context-based pipeline repair slightly reduces matching completeness. (c) F1 Score: Most of the data points are above the equivalent line, indicating that context-based pipeline repair also improves overall matching performance. Experimentally, the wider ellipse around the data points indicates that the matching performance improvement ranking in Phase 2 text preprocessing methods is Stemming (S) > Lemmatisation (L) > Stop Words Removal (R); the matching performance improvement ranking in different stemmers is Lancaster Stemmer (SL) > Porter Stemmer (SP) = Snowball Stemmer (SS); and the matching performance improvement ranking in lemmatisation is Lemmatisation (L) $\approx$ Lemmatisation + POS Tagging (LT).
@@ 6. Text Preprocessing Pipeline & LLMs
In this section, we discuss the pros and cons of LLMs in performing syntactic OM and demonstrate the benefits of integrating the classical text preprocessing pipeline with modern LLMs.
@@ 6.1. Motivation
In modern LLMs, text preprocessing can be performed implicitly within the models. For example, LLMs may understand that “art gallery” and “ArtGallery” are true mappings even if they are in different formats. However, it is questionable whether this behaviour is stable across different LLMs and prompt templates and whether LLMs outperform the classical text processing pipeline. On the other hand, with a strong background knowledge, LLMs may be able to facilitate mapping repair together with the text preprocessing pipeline. For example, LLMs can classify “Bile” and “Bilateral” as false mappings, whereas the classical text preprocessing method stemming cannot.
@@ 6.2. Experiment Setup
    Table 3. Details of selected LLMs (retrieved from  (Qiang et al., 2023)).
    Family Model Size Version GPT gpt-4o N/A gpt-4o-2024-05-13 gpt-4o-mini N/A gpt-4o-mini-2024-07-18 Claude claude-3-sonnet N/A claude-3-sonnet-20240229 claude-3-haiku N/A claude-3-haiku-20240307 Llama llama-3-8b* 4.7 GB Ollama Model ID: 365c0bd3c000 llama-3.1-8b* 4.9 GB Ollama Model ID: 46e0c10c039e Qwen qwen-2-7b* 4.4 GB Ollama Model ID: dd314f039b9d qwen-2.5-7b* 4.7 GB Ollama Model ID: 845dbda0ea48 Gemma gemma-2-9b* 5.4 GB Ollama Model ID: ff02c3702f32 GLM glm-4-9b* 5.5 GB Ollama Model ID: 5b699761eca5 * Open-source LLM (retrieved December 1, 2024).
@@
    We run the experiment on the Largebio Track that contains pairwise alignments between FMA, NCI, and SNOMED. The true mappings (TPs) are retrieved from the reference (R). The false mappings (FPs) are retrieved from the alignment (A) generated by the text preprocessing pipeline but exclude the true mappings (TPs) listed in the reference (R).
@@
    We report our results over a single run. There could be slight differences in the results over multiple trails, but these differences are not significant with respect to the ranking and trend of different LLMs and prompt templates.
@@ 6.3. Detecting True Mappings
LLMs are expected to answer “yes” for TPs. The discovery rate is defined as the number of “yes” answers compared to the total number of TPs. For LLMs, we directly prompt LLMs to determine whether two entities are identical or not. For the text preprocessing pipeline, we employ the pipeline via function calling (a.k.a tool calling). This means that the LLMs act as an interface, and mapping judgement is based on programme logic defined in the text preprocessing method. Two entities are considered to be matched if they have the same results after the text preprocessing.
Figure 10 shows the discovery rate of TPs using LLMs on the Largebio Track. Here we use only Tokenisation (T) and Normalisation (N) to preprocess the text because Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are shown to be less effective from our analysis in Section 4. Although some API-accessed commercial LLMs have shown the capability to detect TPs without a text preprocessing step, this behaviour is non-deterministic across different LLMs and prompt templates. For different LLMs, open-source LLMs perform better than commercial API-accessed LLMs. It could be problematic for some commercial API-accessed LLMs (e.g. claude-3haiku) to find two identical entities if text preprocessing is not applied prior to the comparison. They may consider entities with different formatting that are used in different contexts and therefore have different meanings. For different prompt templates, few-shot (PT2) significantly improves correctness, but self-reflection (PT3) may even hamper performance. The combination of few-shot and self-reflection (PT4) may even perform worse than the simple basic prompt (PT1). We believe that this phenomenon is mainly caused by LLM “overthinking”, a type of LLM hallucinations in which LLMs could over-analyse user input and generate unnecessarily complex details. We also need to note here that the example generated in the few-shot prompt is based on our observation of the existing TPs. Finding a good example in real-world applications is not easy because ontology entities can have different formatting by nature. On the other hand, using the classical text preprocessing pipeline via function calling could help LLMs produce a stable result with a high discovery rate (almost 100% from our observation) across different models and prompt templates.

### v5

@@ 4.2.3. Different Stemming Methods
Figure 6 compares different stemming methods: Porter Stemmer (SP), Snowball Stemmer (SS), and Lancaster Stemmer (SL). For SP and SS, all data points located on the equivalence line indicate that there is no difference between SP and SS. For SP/SS and SL, most data points above the equivalence line in precision and F1 score indicate that SP/SS is more effective than SL. In the violin plots: (a) Precision: The median number of SP and SS is greater than that of SL, and there is no difference between SP and SS. The shape of the distribution is identical. (b) Recall: The median number and the shape of the distribution are identical in SP, SS, and SL. (c) F1 Score: The median number of SP and SS is greater than that of SL, and there is no difference between SP and SS. The shape of the distribution is identical in SP, SS, and SL.
@@ 4.2.4. Lemmatisation with/without POS Tagging
Figure 7 compares Lemmatisation (L) and Lemmatisation + POS Tagging (LT). All data points located on the equivalence line indicate that using POS Tagging in Lemmatisation does not help syntactic OM. In the violin plots, for each of (a) Precision, (b) Recall, and (c) F1 Score, the median number and the shape of the distribution are identical for L and LT.
@@ 4.3. Implications
    Phase 1 text preprocessing methods: We observe that matching performance usually increases with each Phase 1 text preprocessing method. However, Phase 1 methods could also benefit from customisation in OM tasks. Some traditional methods originating from natural language processing (NLP) are not useful for OM and need to be adjusted. For example, sentence segmentation (i.e. splitting long text into sentences) is not applicable for ontology entities because they are generally short text fragments and do not need such operations. Word tokenisation (i.e. breaking text into single words) could be rewritten to detect the abbreviations that are common practice in ontology concept names and to tackle the conflicting use of Pascal case and snake case.
@@
      Stop Words Removal (R): Logical expressions “and”, “or”, and “not” are stop words in English and usually carry little useful information, whereas these words express logical operations in ontology entities and may carry important semantics. Some single stop word characters (e.g. “d”, “i”, “m”, “o”, “s”, “t”, and “y”) should be removed from the stop word set because contractions (e.g. I’d, I’m, o’clock, it’s, isn’t, y’know) are unlikely to appear in ontology entities, and yet those single characters are likely to carry important meaning if they do occur (e.g. “Chromosome_Y”).
@@
      Stemming (S) vs. Lemmatisation (L): While L tends to avoid generating FPs, it may also miss some implicit TPs. On the other hand, S is more eager to find TPs, but eagerness can also lead to more FPs being found. Based on our experiments with no post hoc correction applied, L is better than S.
@@
      Different stemming methods: Porter Stemmer (PS) and Snowball Stemmer (SS) have been found to perform better than Lancaster Stemmer (LS), and we cannot see a significant difference between Porter and Snowball. Although the Lancaster Stemmer is more eager to detect more TPs, it does not lead to performance improvement as more FPs are also matched.
@@ 5.1. Causal Analysis
    Phase 1 text preprocessing methods (Tokenisation and Normalisation) do not change the text meaning. For example, isReviewing is equivalent to is_reviewing. This means that applying Phase 1 methods only help to detect TPs, while the number of FPs remains unchanged. For this reason, Phase 1 methods always have a positive effect on precision, recall, and overall F1 score.
@@
    Phase 2 text preprocessing methods (Stop Words Removal and Stemming/Lemmatisation) are actually relaxations of matching rules in OM tasks. Moving through the text preprocessing pipeline tends to detect more TPs and FPs in the derived alignment (A). For example, isReviewing and isReviewedBy may be object properties with distinctly different meanings, but removing the common stop words “is” and “by”, and using Stemming/Lemmatisation to retrieve the same root word “Review”, could cause a false match. For this reason, recall is always increasing, but precision and the overall F1 score are less reliable.
@@
    If we define $\Delta TP$ and $\Delta FP$ as the increase in TP and FP for a preprocessing method, then the condition to increase precision and F1 score is:
@@
    In our experiments, we actually observe a reduction in precision and overall F1 score. This means that Phase 2 methods produce more FPs than TPs, and this proportion is less than the original number of $TP/FP$. So performance does not improve, unless the benefit of each TP is considered more valuable than the disbenefit of each FP. This could apply, for example, if we are expecting a post hoc correction phase in which removing FPs is considered to be an easier human task than adding missing TPs. This idea inspires our novel pipeline repair method.
@@ 5.2. Repair Method
While Phase 2 text preprocessing methods have been shown to be less effective in syntactic OM by introducing FPs, we propose a pipeline repair approach to differentiate FPs from TPs. As illustrated in Figure 8: (1) Phase 1 methods shift Alignment (A) towards Reference (R). The number of TPs increases, while the number of FPs decreases. (2) Phase 2 methods expand Alignment (A). The number of TPs increases, but the number of FPs also increases. (3) Our context-based pipeline repair aims to collapse Alignment (A). The number of FPs decreases significantly, with a slight decrease in TPs.
@@
One critical step in our approach is to assemble a set of reserved words that may cause FPs in text preprocessing, and to exclude these words from text preprocessing. The selection criteria are based on two widely agreed assumptions: (1) there are no duplicate entities within a single ontology; and (2) ontologies that represent the same domain knowledge tend to use similar terminologies. We call our approach context-based here because pairs of words may have the same or different meanings in different contexts. Based on these two assumptions, we propose a simple Algorithm 1 to assemble the reserved word set. For both $O_{s}$ and $O_{t}$, the algorithm first iterates on all pairs of entities $e_{i},e_{j}$ in the ontology, respectively. For a specific text preprocessing method $f(\cdot)$, if $f(e_{i})=f(e_{j})$, we retrieve the different words between $e_{i}$ and $e_{j}$ and store them in the reserved word set. If a word appears in the reserved set, the text preprocessing pipeline skips the operation for this word. To simplify the reserved word set, we remove the words where $f(w)=w$ because their occurrence in the reserved word set has no effect on mapping results.
@@
  /* Phase 2: Find immutable words in $Reserved\_Word\_Set$/
@@
An example of generating and using a simple reserved word set is illustrated below. (1) The two object properties was_a_member_of and has_members from a single ontology have the same result “member” via traditional text preprocessing. Because there are no duplicate entities within a single ontology, we use a reserved word set in our proposed pipeline repair approach to determine that they are distinguishing entities. The initial step (i.e. Phase 1 of Algorithm 1) is to add [“was”, “a”, “member”, “of”, “has”, “members”] to the reserved word set so that these two object properties would not have the same text preprocessing results. was_a_member_of preprocessed with skipping the reserved word set is “was a member of”, while has_members preprocessed with the reserved word set is “has members”. The revision step (i.e. Phase 2 of Algorithm 1) is to check whether there are immutable words in the reserved set. We can observe that the word “member” is immutable because it is the same before and after text preprocessing. Removing this word from the reserved word set still makes the two object properties different. Therefore, the final reserved word set is [“was”, “a”, “of”, “has”, “members”]. (2) The generated reserved word set can be used to repair false mappings between entities within the same domain context but coming from different ontologies. For example, we expect that the two object properties has_a_steering_committee and was_a_steering_committee_of are non-identical. While a false mapping may occur when they both have the same result “steer committe” after the traditional text preprocessing, using the reserved word set can repair this false mapping. As the words “has”, “a”, “was”, and “of” are listed as reserved words, these two named properties are processed as “has a steer committe” and “was a steer committe of”, respectively.
@@ 5.3. Evaluation
We apply our context-based pipeline repair approach to the same OAEI datasets as above. Figure 9 compares with and without context-based pipeline repair in 8 tracks with 49 distinct alignments. The text preprocessing pipeline methods implemented include all Phase 2 text preprocessing methods: Stop Words Removal (R), Porter Stemmer (SP), Snowball Stemmer (SS), Lancaster Stemmer (SL), Lemmatisation (L), and Lemmatisation + POS Tagging (LT).
@@
  1. (1)
    Precision: Most data points are above the equivalence line, indicating that context-based pipeline repair significantly improves matching correctness. Recall: Most data points are located on the equivalence line, and only a few data points are below the equivalence line, indicating that context-based pipeline repair slightly reduces matching completeness. F1 Score: Most of the data points are above the equivalence line, indicating that context-based pipeline repair also improves overall matching performance.
  2. (2)
    Experimentally, the wider ellipse around the data points indicates that the matching performance improvement ranking in Phase 2 text preprocessing methods is Stemming (S) > Lemmatisation (L) > Stop Words Removal (R); the matching performance improvement ranking in different stemmers is Lancaster Stemmer (SL) > Porter Stemmer (SP) = Snowball Stemmer (SS); and the matching performance improvement ranking in lemmatisation is Lemmatisation (L) $\approx$ Lemmatisation + POS Tagging (LT).
@@ 6. Text Preprocessing Pipeline & LLMs
In this section, we discuss the pros and cons of LLMs in performing syntactic OM and demonstrate the benefits of integrating the classic text preprocessing pipeline with modern LLMs.
@@ 6.1. Motivation
In modern LLMs, text preprocessing can be performed implicitly within the models. For example, LLMs may understand that “art gallery” and “ArtGallery” are true mappings even if they are in different formats. However, it is questionable whether this behaviour is stable across different LLMs and prompt templates and whether LLMs outperform the classic text processing pipeline. On the other hand, with a strong background knowledge, LLMs may be able to facilitate mapping repair together with the text preprocessing pipeline. For example, LLMs can classify “Bile” and “Bilateral” as false mappings, whereas the classic text preprocessing method stemming cannot.
@@ 6.2. Experiment Setup
    Table 3. Details of selected LLMs (retrieved from (Qiang et al., 2023)). API-accessed models are shown as triangles ($\blacktriangle$) and open-source models as circles ($\bullet$).
    Family Model Version Type GPT gpt-4o gpt-4o-2024-05-13 $\blacktriangle$ gpt-4o-mini gpt-4o-mini-2024-07-18 $\blacktriangle$ Claude claude-3-sonnet claude-3-sonnet-20240229 $\blacktriangle$ claude-3-haiku claude-3-haiku-20240307 $\blacktriangle$ Llama llama-3-8b Ollama Model ID: 365c0bd3c000 $\bullet$ llama-3.1-8b Ollama Model ID: 46e0c10c039e $\bullet$ Qwen qwen-2-7b Ollama Model ID: dd314f039b9d $\bullet$ qwen-2.5-7b Ollama Model ID: 845dbda0ea48 $\bullet$ Gemma gemma-2-9b Ollama Model ID: ff02c3702f32 $\bullet$ GLM glm-4-9b Ollama Model ID: 5b699761eca5 $\bullet$
@@
    We run the experiment on the Largebio Track that contains pairwise alignments between FMA, NCI, and SNOMED. The true mappings (TPs) are retrieved from the reference (R) where the mappings can be found by syntactic OM. The false mappings (FPs) are retrieved from the alignment (A) generated by the classic text preprocessing pipeline but exclude TPs above.
@@
    We report our results over a single run. There could be slight differences in the results over multiple trials, but these differences are not significant with respect to the ranking and trend of different LLMs and prompt templates.
@@ 6.3. Detecting True Mappings
LLMs are expected to answer “yes” for TPs. The discovery rate is defined as the number of “yes” answers compared to the total number of TPs. For LLMs, we directly prompt LLMs to determine whether two entities are identical or not. For the text preprocessing pipeline, we employ the pipeline via function calling (OpenAI, 2023) (a.k.a. tool calling) in LLMs. This means that the LLMs act as an interface, and mapping judgement is based on programme logic defined in the text preprocessing method. Two entities are considered to be matched if they have the same results after the text preprocessing.
Figure 10 shows the discovery rate of TPs using LLMs on the Largebio Track. Here we use only Tokenisation (T) and Normalisation (N) to preprocess the text because Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are shown to be less effective in our analysis in Section 4. Although some API-accessed commercial LLMs have shown the capability to detect TPs without a text preprocessing step, this behaviour varies across different LLMs and prompt templates. For different LLMs, open-source LLMs perform better than commercial API-accessed LLMs. It could be problematic for some commercial API-accessed LLMs (e.g. claude-3-haiku) to find two identical entities if text preprocessing is not applied prior to the comparison. They may consider entities with different formatting that are used in different contexts and therefore have different meanings. For different prompt templates, few-shot (PT2) significantly improves correctness, but self-reflection (PT3) may even hamper performance. The combination of few-shot and self-reflection (PT4) may even perform worse than the simple basic prompt (PT1). We believe that this phenomenon is mainly caused by “overthinking”, a type of LLM hallucination in which LLMs could over-analyse user input and generate unnecessarily complex details (Chen et al., 2025). We also need to note here that the example generated in the few-shot prompt is based on our observation of the existing TPs. Finding a good example in real-world applications is not easy because ontology entities can have different formatting by nature. On the other hand, using the classic text preprocessing pipeline via function calling could help LLMs produce a stable result with a high discovery rate (almost 100% from our observation) across different models and prompt templates.